OpenStudy (anonymous):

Can someone help me find the partial sum of:

OpenStudy (anonymous):

\[\frac{ 1 }{n(n+1)}\] that's the term of the series

OpenStudy (anonymous):

The series begins at n=1

OpenStudy (anonymous):

and end up at... ?

OpenStudy (anonymous):

Maybe at "n" @Diogo

OpenStudy (anonymous):

then just put a sigma from n=1 to +oo before

OpenStudy (anonymous):

It ends up at infinity

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ 1 }{ n (n+1) }\]

OpenStudy (amistre64):

partial sums may form patterns that can be defined explicitly; otherwise it might be best to do a partial decomp and try to do a term by term approach

OpenStudy (anonymous):

isn't there an other way?

OpenStudy (amistre64):

im sure there are plenty of methods out there ...

OpenStudy (amistre64):

\[\frac12=\frac12\\ \frac12+\frac16=\frac{8}{12}\\ \frac23+\frac1{12}=\frac{27}{36}\\\] etc ..

OpenStudy (amistre64):

see if the bottom and top can be defined as sequences

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ 1 }{ n(n+1) } = 1- \frac{ 1 }{\infty } = 1 - 0 =\]

OpenStudy (amistre64):

\[S_n=\frac{n}{n+1}\]maybe

OpenStudy (anonymous):

Actually I don't understand what Diogo wrote...

OpenStudy (amistre64):

Diogo was interpreting this as the ultimate partial sum, the TOTAL sum as n goes to infinity

OpenStudy (amistre64):

the partial sums are the intermediary sums from 1 to n, for a given integer n

OpenStudy (anonymous):

How does he get to the final value then?

OpenStudy (amistre64):

he has his ways :)

OpenStudy (amistre64):

but the question, as far as i can tell, is not asking for the SUM of the series as a whole, but rather as an explicit rule in terms of n. as n approaches infinity, the rule simplifies to 1

OpenStudy (amistre64):

30 minutes to a final .. ive got to get a little refresher studying in

OpenStudy (anonymous):

I think I got it! so you make the partial sum of the first few numbers and then you try to recognize a pattern. thanks

OpenStudy (anonymous):

\[ \frac{ 1 }{ n(n+1) } = \frac{ 1+n -n }{ n(n+1) } = \frac{ n+1 }{ n(n+1) } - \frac{ n }{ n(n+1) } = \frac{ 1 }{ n } - \frac{ 1 }{ (n+1) } \] for n = 1 you have 1-0.5 for n = 2 you have 0.5-0.3333 for n=3 you have 0.3333-0.25 If you sum all you'll have 1-0.5 + 0.5 - 0.3333 +0.3333-0.25 +0.25 + .... as you can see all the terms cancel each other so the final answer is just 1

OpenStudy (anonymous):

ok thanks